# Multifidelity Simulation-based Inference for Computationally Expensive Simulators

### v3

@@ 3.1.2 Sequential training
In addition to learning amortized posterior estimates with NPE, our approach naturally extends to sequential training schemes when estimating the non-amortized posterior $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x_{o}})$. Rather than sampling model parameters from the prior, sequential methods introduce an active learning scheme that iteratively refines the posterior estimate for a specific observation $\boldsymbol{x_{o}}$. These methods – known as Sequential Neural Posterior Estimation (Papamakarios & Murray, 2016; Lueckmann et al., 2017) – have shown increased simulation efficiency when compared to NPE (Lueckmann et al., 2021). However, applying these methods with flexible neural density estimators requires a modified loss that suffers from instabilities in training and posterior leakage (Greenberg et al., 2019). Truncated Sequential Neural Posterior Estimation (TSNPE) mitigates these issues by sampling from a truncated prior distribution that covers the support of the posterior. This leads to a simplified loss function and increased training stability, while retaining performance (Deistler et al., 2022).
@@ 3.1.3 Acquisition function
To further enhance the efficiency of our sequential algorithm, we explore the use of acquisition functions to supplement our round-wise samples from the TNSPE proposal: we generate simulations for round $i$ with a set of parameters $\boldsymbol{\theta}^{(i)}=\ \{\boldsymbol{\theta}^{(i)}_{\mathrm{prop}}\cup\boldsymbol{\theta}^{(i)}_{\mathrm{active}}\}$ where $\boldsymbol{\theta}^{(i)}_{\mathrm{prop}}$ are samples from the proposal distribution at round $i$, and $\boldsymbol{\theta}^{(i)}_{\mathrm{active}}$ are the top $\mathcal{B}$ values according to an acquisition function. We refer to this algorithm as A-MF-TSNPE (full description in Appendix M.2). Following Järvenpää et al. (2019); Lueckmann et al. (2019), we select an acquisition function that targets the variance of the posterior estimate with respect to the uncertainty in the learned parameters $\phi|\mathcal{D}$.
@@
We realize this as the sample variance across an ensemble of neural density estimators trained independently on the same dataset $\mathcal{D}$, as done in Lueckmann et al. (2019). Alternatively, one could estimate $\phi|\mathcal{D}$ using other methods of uncertainty quantification, such as Monte Carlo dropout (Griesemer et al., 2024). For details on the proposal design of A-MF-TSNPE, see Appendix M.2.
@@ 3.1.4 Evaluation metrics
We evaluate the accuracy of posterior distributions in cases where the ground-truth posterior is known with the Classifier-2-Sample Test (C2ST) and the Maximum Mean Discrepancy (MMD)(Friedman, 2004; Lopez-Paz & Oquab, 2017; Gretton et al., 2012; Lueckmann et al., 2021; Peyré & Cuturi, 2017). C2ST is commonly used in SBI, as it is easy to apply and interpret: a value close to 0.5 means that a classifier cannot effectively distinguish the two distributions, implying the posterior estimate is close to the ground-truth posterior. A value close to 1 means that the classifier can distinguish the distributions very well, indicating a poor posterior estimation. C2ST is rarely applicable in practical SBI settings, since it requires samples from the true posterior (e.g., Sec. 4.1).
@@
The average Negative Log probability of the True Parameters (NLTP; $\mathbb{E}[\log q(\boldsymbol{\theta_{o}}|\boldsymbol{x_{o}})]$ ) has been extensively used in the SBI literature for problems where the true posterior is unknown (Greenberg et al., 2019; Papamakarios & Murray, 2016; Durkan et al., 2020; Hermans et al., 2020). In the limit of a large number of pairs $(\boldsymbol{\theta_{o}},\boldsymbol{x_{o}})$, the average over the log probability of each pair $(\boldsymbol{\theta_{o}},\boldsymbol{x_{o}})$ approaches the expected KL divergence between the estimated and the true posterior (up to a term that is independent of the estimated posterior), as shown in (Lueckmann et al., 2021). In addition, we report the Normalized Root Mean Square Error (NRMSE), which quantifies the deviation of posterior samples from the true parameters on a scale-invariant axis. NRMSE values closer to 0 indicate better predictive performance.
@@ 4 Results
We evaluate the performance of our multifidelity approach to NPE and TSNPE on six tasks involving various types of observations (e.g., time series, images, neural spiking). We start with four benchmarking tasks, followed by two challenging neuroscience problems with computationally expensive simulators and for which no likelihood is available: a multicompartmental neuron model and a neural network model with synaptic plasticity. We also provide a comparison to MF-ABC (Sec. F.1.1, E.3).
@@ 4.1 Benchmarking tasks
We first evaluated MF-(TS)NPE on four benchmarking tasks: SIR, SLCP, OUprocess, and Gaussian Blob. SIR and SLCP are established SBI benchmarks (Lueckmann et al., 2021), OUprocess is a new multifidelity task with tractable likelihood (Kou et al., 2012), and Gaussian Blob is a high-dimensional image task (Lueckmann et al., 2019) (details in Appendix E). These tasks were chosen to systematically investigate various task properties that might impact the performance of transfer learning in a multifidelity setting: differing parameter dimensionality between the low- and high-fidelity models, partly observed dynamics, differing simulator types between the low- and high-fidelity models, and high-dimensional observations. Furthermore, these multifidelity tasks are not trivial in the sense that the low and high-fidelity simulators lead to different posteriors (Appendix J). Note that we do not evaluate the total cost of low- and high-fidelity simulations in these tasks, but defer this analysis to the two complex neuroscience tasks (Appendix K).
@@
To evaluate MF-NPE, we compared the estimated densities to the respective reference posterior, estimated from the exact likelihood with Rejection Sampling (Martino et al., 2018) (OU process; closed-form of the likelihood in Sec. E.1), and using Sampling and Importance Resampling (RUBIN, 1988) to obtain a set of 10k proposal samples (SLCP, SIR), similar to Lueckmann et al. (2021). We quantified the performance with C2ST and MMD over 10 observations (30 observations for the OU process) and 10 network initializations per observation. GaussianBlob uses a CNN embedding and was evaluated with NRMSE and NLTP since no closed-form likelihood is available (Fig. 10).
@@
Figure 2: (A)-MF-(TS)NPE outperforms NPE and TSNPE in simulation-efficiency. C2ST and MMD averaged over 10 network initializations with means and $95\%$ confidence intervals. MF-NPE4 and MF-NPE5 are pretrained on $10^{4}$ and $10^{5}$ low-fidelity simulations, respectively. Results for the GaussianBlob task in Fig. 10; variations on the OU task and comparisons to MF-ABC in Fig. 8.
@@
Across experiments, we observed a consistent performance increase with MF-NPE compared to NPE, and (A)-MF-TSNPE compared to TSNPE, especially in low simulation budgets from the high-fidelity model (50-$10^{3}$ simulations) (Fig. 2; Gaussian Blob in Fig. 10). In addition, we found that having a higher number of low-fidelity samples improved performance, reinforcing that low-fidelity simulations were indeed advantageous for pre-training the neural density estimator for the downstream task. Note that for the OU and SLCP tasks, we did not observe a substantial increase in MF-NPE performance between the settings with $10^{4}$ and $10^{5}$ low-fidelity samples, suggesting an upper bound regarding pre-training efficacy. We also compared MF-NPE with MF-ABC, an ABC-based method for multifidelity SBI (Prescott & Baker, 2020), and observed that MF-NPE has a substantially higher performance (Appendix F.1.1). This is consistent with previous findings indicating the superior performance of NPE with respect to rejection ABC and SMC-ABC, where it is not uncommon to require orders of magnitude more simulations to obtain reliable posterior approximations (Lueckmann et al., 2021; Frazier et al., 2024). However, a more extensive hyperparameter search could potentially lead to substantial improvements in MF-ABC performance.
@@
As described in Sec. 3, we enhanced the sequential algorithm TSNPE (Deistler et al., 2022) with a first round of MF-NPE, and designated this approach as MF-TSNPE. We found that MF-TSNPE (details in Appendix M.1) performs better than TSNPE, especially in regimes with a low budget of high-fidelity simulations. Compared to MF-TSNPE, A-MF-TSNPE improved inference in the OU process, but did not show significant improvements in the SLCP and SIR tasks.
@@
Finally, we assessed the contribution of transfer learning to the overall performance in a setting where the low- and high-fidelity models have a different number of parameters, in the context of the OUprocess task (Appendix E.3). We expected that adding parameters to the high-fidelity model that are absent in the low-fidelity model would increase the inference complexity for MF-NPE, and indeed observed a performance decrease in MF-NPE, although MF-NPE still performed better than NPE and MF-ABC (see Appendix E.3). Overall, the results suggest that MF-NPE and MF-TSNPE can yield substantial performance gains compared to NPE, TSNPE, and MF-ABC.
@@ 4.2 Multicompartmental neuron model
The voltage response of a morphologically-detailed neuron to an input current is typically modeled with a multicompartment model wherein the voltage dynamics of each compartment are based on the Hodgkin-Huxley equations (Hodgkin & Huxley, 1952). The higher the number of compartments of the model, the more accurate the model is, but the higher the simulation cost.
@@
To simulate the neuron models, we used Jaxley, a Python toolbox for efficiently simulating multicompartment single neurons with biophysical detail (Deistler et al., 2024). In this setting, the simulation time for the high-fidelity model is approximately 4 times higher than that of the low-fidelity model. We characterized the neural response with four summary statistics that have been commonly used when fitting biophysical models of single neurons to empirical data: spike count, mean resting potential, standard deviation of the resting potential, and voltage mean (Gonçalves et al., 2020; Gao et al., 2023). Performances were evaluated with NLTP and NRMSE on $10^{3}$ pairs of $\theta_{o}$ and respective simulation outputs $x_{o}$, averaged over 10 random network initializations (Sec. 3.1.4).
MF-(TS)NPE showed higher performance than NPE, in particular with larger low-fidelity simulation budgets (Fig. 3B; Fig. G.1), despite the right-skewed posterior distribution of the low-fidelity model (Fig. 19). Furthermore, MF-NPE posterior predictives closely matched the empirical data, in contrast with NPE, even when NPE was trained on a higher number of high-fidelity simulations (Appendix G). In addition, MF-(TS)NPE achieved comparable performance with a total computational cost 4-5 times lower than standard NPE (Appendix K). Finally, simulation-based calibration tests suggest that both MF-NPE and NPE estimates were relatively well calibrated (Fig. 3C) (Talts et al., 2020).
(A)-MF-TSNPE pre-trained on $10^{4}$ low-fidelity samples outperforms MF-NPE trained on $10^{5}$ samples. However, A-MF-TSNPE performance comes at the cost of training time due to the use of an ensemble of density estimators (Appendix K). This additional training burden is only justified when the simulation cost is substantially higher than the training cost.
Figure 3: MF-NPE improves inference on neuroscience tasks. (A) Thick-tufted layer 5 pyramidal cell from the neocortex. (B) Performance evaluation with NLTP (same naming convention as in Fig. 2). Amortized methods are averaged over 10 network initializations; non-amortized trained once per 100 observations. Similar results were obtained with NRMSE (Appendix G.1). MF-NPE, and especially its sequential variants, are orders of magnitude more simulation-efficient than NPE. (C) Simulation-based calibration for NPE and MF-NPE trained on $10^{3}$ high-fidelity samples. Posterior samples and predictives are in Appendix G. (D) Schematic of the low and high-fidelity models of a spiking network. (E) Performance of NPE and MF-NPE evaluated on 10000 true observations with NLTP: averages over 10 network initializations, and 95% confidence intervals. (F) Proportion of posterior samples within the target firing rate bounds. MF-NPE produces a higher fraction of parameter sets within the bounds than NPE.
@@ 4.3 Recurrent spiking network
The high-fidelity simulator consisted of a recurrent network of 4096 excitatory ($E$) and 1024 inhibitory ($I$) leaky integrate-and-fire neurons connected with conductance-based synapses (Fig. 3D). Each synapse type in this network ($E$-to-$E$, $E$-to-II, II-to-$E$, II-to-II) was plastic with an unsupervised local learning rule. For each synapse type, 6 parameters governed how the recent pre- and post-synaptic activity were used to update the synapse, for a total of 24 free parameters across all 4 synapse types (Confavreux et al., 2023). The networks were simulated using Auryn, a C++ simulator (Zenke & Gerstner, 2014) (details in Appendix H).
@@
Summary statistics of the low- and high-fidelity models were the average firing rates of the excitatory and inhibitory neurons at steady state (after 2 minutes of simulation in the high-fidelity model). Plastic networks were considered plausible if the firing rates were between 1 and 50Hz (Dayan & Abbott, 2001; Confavreux et al., 2023).
In this task, the low-fidelity model focuses solely on the $E$-to-$E$ and $E$-to-II rules from the high-fidelity model, thereby having 12 out of the 24 parameters of the high-fidelity model. This setup allows us to demonstrate the performance of MF-NPE on problems with different parameter spaces, highlighting MF-NPE’s flexibility and advantages. We found that MF-NPE has better performance than NPE in terms of NLTP (Fig. 3E). Furthermore, MF-NPE leads to an increase of almost $30\%$ in the proportion of posterior samples within the target firing rate bounds (Fig. 3F), reinforcing that MF-NPE is a practical and effective method for SBI of costly real-world simulators.

### v4

@@ 3.1.2 Sequential training
In addition to learning amortized posterior estimates with NPE, our approach naturally extends to sequential training schemes when estimating the non-amortized posterior $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x_{o}})$. Rather than sampling model parameters from the prior, sequential methods introduce an active learning scheme that iteratively refines the posterior estimate for a specific observation $\boldsymbol{x_{o}}$. These methods – known as Sequential Neural Posterior Estimation (Papamakarios and Murray, 2016; Lueckmann et al., 2017) – have shown increased simulation efficiency when compared to NPE (Lueckmann et al., 2021). However, applying these methods with flexible neural density estimators requires a modified loss that suffers from instabilities in training and posterior leakage (Greenberg et al., 2019). Truncated Sequential Neural Posterior Estimation (TSNPE) mitigates these issues by sampling from a truncated prior distribution that covers the support of the posterior. This leads to a simplified loss function and increased training stability, while retaining performance (Deistler et al., 2022).
@@ 3.1.3 Acquisition function
To further enhance the efficiency of our sequential algorithm, we explore the use of acquisition functions to supplement our round-wise samples from the TSNPE proposal: we generate simulations for round $i$ with a set of parameters $\boldsymbol{\theta}^{(i)}=\ \{\boldsymbol{\theta}^{(i)}_{\mathrm{prop}}\cup\boldsymbol{\theta}^{(i)}_{\mathrm{active}}\}$ where $\boldsymbol{\theta}^{(i)}_{\mathrm{prop}}$ are samples from the proposal distribution at round $i$, and $\boldsymbol{\theta}^{(i)}_{\mathrm{active}}$ are the top $\mathcal{B}$ values according to an acquisition function. We refer to this algorithm as MF-TSNPE-AF (full description in Appendix M.2). Following Järvenpää et al. (2019); Lueckmann et al. (2019), we select an acquisition function that targets the variance of the posterior estimate with respect to the epistemic uncertainty in the learned parameters $\phi|\mathcal{D}$.
@@
We realize this as the sample variance across an ensemble of neural density estimators trained independently on the same dataset $\mathcal{D}$, as done in Lueckmann et al. (2019). Note that we use epistemic uncertainty to guide high-fidelity simulation selection within the simulator’s domain rather than out-of-distribution samples. For details on the proposal design of MF-TSNPE-AF, see Appendix M.2.
@@ 3.1.4 Evaluation metrics
We evaluate the accuracy of posterior distributions in cases where the ground-truth posterior is known with the Classifier-2-Sample Test (C2ST) and the Maximum Mean Discrepancy (MMD)(Friedman, 2004; Lopez-Paz and Oquab, 2017; Gretton et al., 2012; Lueckmann et al., 2021; Peyré and Cuturi, 2017). C2ST is commonly used in SBI, as it is easy to apply and interpret: a value close to 0.5 means that a classifier cannot effectively distinguish the two distributions, implying the posterior estimate is close to the ground-truth posterior. A value close to 1 means that the classifier can distinguish the distributions very well, indicating a poor posterior estimation. C2ST is rarely applicable in practical SBI settings, since it requires samples from the true posterior (e.g., Sec. 4.1).
@@
The average Negative Log probability of the True Parameters (NLTP; $\mathbb{E}[\log q(\boldsymbol{\theta_{o}}|\boldsymbol{x_{o}})]$ ) has been extensively used in the SBI literature for problems where the true posterior is unknown (Greenberg et al., 2019; Papamakarios and Murray, 2016; Durkan et al., 2020; Hermans et al., 2020). In the limit of a large number of pairs $(\boldsymbol{\theta_{o}},\boldsymbol{x_{o}})$, the average over the log probability of each pair $(\boldsymbol{\theta_{o}},\boldsymbol{x_{o}})$ approaches the expected KL divergence between the estimated and the true posterior (up to a term that is independent of the estimated posterior), as shown in (Lueckmann et al., 2021). In addition, we report the Normalized Root Mean Square Error (NRMSE), which quantifies the deviation of posterior samples from the true parameters on a scale-invariant axis. NRMSE values closer to 0 indicate better predictive performance.
@@ 4 Results
We evaluate the performance of our multifidelity approach to NPE and TSNPE on six tasks involving various types of observations (e.g., time series, images, neural spiking). We start with four benchmarking tasks, followed by two challenging neuroscience problems with computationally expensive simulators and for which no likelihood is available: a multicompartmental neuron model and a neural network model with synaptic plasticity. We also provide a comparison to MF-ABC (Sec. E.1.1, D.3). In Sec. 4.4, we provide a discussion about the effectiveness of transfer learning in MF-NPE.
@@ 4.1 Benchmarking tasks
We first evaluated MF-(TS)NPE on four benchmarking tasks: SIR, SLCP, OUprocess, and Gaussian Blob. SIR and SLCP are established SBI benchmarks (Lueckmann et al., 2021), OUprocess is a new multifidelity task with tractable likelihood (Kou et al., 2012), and Gaussian Blob is a high-dimensional image task (Lueckmann et al., 2019) (details in Appendix D). These tasks were chosen to systematically investigate various task properties that might impact the performance of transfer learning in a multifidelity setting: differing parameter dimensionality between the low- and high-fidelity models, partly observed dynamics, differing simulator types between the low- and high-fidelity models, and high-dimensional observations. Furthermore, these multifidelity tasks are not trivial in the sense that the low and high-fidelity simulators lead to different posteriors (Appendix I). Note that we do not evaluate the total cost of low- and high-fidelity simulations in these tasks, but defer this analysis to the two complex neuroscience tasks (Appendix J).
@@
To evaluate MF-NPE, we compared the estimated densities to the respective reference posterior, estimated from the exact likelihood with Rejection Sampling (Martino et al., 2018) (OU process; closed-form of the likelihood in Sec. D.1), and using Sampling and Importance Resampling (RUBIN, 1988) to obtain a set of 10k proposal samples (SLCP, SIR), similar to Lueckmann et al. (2021). We quantified the performance with C2ST and MMD over 10 observations (30 observations for the OU process) and 10 network initializations per observation. GaussianBlob required a CNN embedding and was evaluated with NRMSE and NLTP since no closed-form likelihood is available (Fig. 11).
@@
Figure 2: C2ST and MMD averaged over 10 network initializations with means and $95\%$ confidence intervals. MF-NPE4 and MF-NPE5 are pretrained on $10^{4}$ and $10^{5}$ low-fidelity simulations, respectively. Results for the GaussianBlob task in Fig. 11; variations on the OU task and comparisons to MF-ABC in Fig. 8.
@@
Across four benchmarking tasks, we observed a consistent performance increase with MF-NPE compared to NPE, and MF-TSNPE(-AF) compared to TSNPE, especially in low simulation budgets from the high-fidelity model (50-$10^{3}$ simulations) (Fig. 2; Gaussian Blob in Fig. 11). In addition, we found that having a higher number of low-fidelity samples improved performance, reinforcing that low-fidelity simulations were indeed advantageous for pre-training the neural density estimator for the downstream task. Note that for the OU and SLCP tasks, we did not observe a substantial increase in MF-NPE performance between the settings with $10^{4}$ and $10^{5}$ low-fidelity samples, suggesting an upper bound regarding pre-training efficacy. We also compared MF-NPE with MF-ABC, an ABC-based method for multifidelity SBI (Prescott and Baker, 2020), and observed that MF-NPE has a substantially higher performance (Appendix E.1.1). This is consistent with previous findings indicating the superior performance of NPE with respect to rejection ABC and SMC-ABC, where it is not uncommon to require orders of magnitude more simulations to obtain reliable posterior approximations (Lueckmann et al., 2021; Frazier et al., 2024). However, a more extensive hyperparameter search could potentially lead to substantial improvements in MF-ABC performance.
@@
As described in Sec. 3, we enhanced the sequential algorithm TSNPE (Deistler et al., 2022) with a first round of MF-NPE, and designated this approach as MF-TSNPE. We found that MF-TSNPE (details in Appendix M.1) performs better than TSNPE, especially in regimes with a low budget of high-fidelity simulations. Compared to MF-TSNPE, MF-TSNPE-AF improved inference in the OU process, but did not show significant improvements in the SLCP and SIR tasks.
@@
Finally, we assessed the contribution of transfer learning to the overall performance in a setting where the low- and high-fidelity models have a different number of parameters, in the context of the OUprocess task (Appendix D.3). We expected that adding parameters to the high-fidelity model that are absent in the low-fidelity model would increase the inference complexity for MF-NPE, and indeed observed a performance decrease in MF-NPE, although MF-NPE still performed better than NPE and MF-ABC (see Appendix D.3). We note that MF-NPE also outperformed NPE when the low-fidelity model had more parameters than the high-fidelity model (see Appendix D.4). Overall, the results suggest that MF-NPE and MF-TSNPE can yield substantial performance gains compared to NPE, TSNPE, and MF-ABC.
@@ 4.2 Multicompartmental neuron model
The voltage response of a morphologically-detailed neuron to an input current is typically modeled with a multicompartment model wherein the voltage dynamics of each compartment are based on the Hodgkin-Huxley equations (Hodgkin and Huxley, 1952). The higher the number of compartments of the model, the more accurate the model is, but the higher the simulation cost.
@@
To simulate the neuron models, we used Jaxley, a Python toolbox for efficiently simulating multicompartment single neurons with biophysical detail (Deistler et al., 2025). In this setting, the simulation time for the high-fidelity model is approximately 4 times higher than that of the low-fidelity model. We characterized the neural response with four summary statistics that have been commonly used when fitting biophysical models of single neurons to empirical data: spike count, mean resting potential, standard deviation of the resting potential, and voltage mean (Gonçalves et al., 2020; Gao et al., 2023). Performances were evaluated with NLTP and NRMSE on $10^{3}$ pairs of $\theta_{o}$ and respective simulation outputs $x_{o}$, averaged over 10 random network initializations (Sec. 3.1.4).
MF-(TS)NPE showed higher performance than NPE, in particular with larger low-fidelity simulation budgets (Fig. 3B; Fig. F.1), despite the right-skewed posterior distribution of the low-fidelity model (Fig. 21). Furthermore, MF-NPE posterior predictives closely matched the empirical data, in contrast with NPE, even when NPE was trained on a higher number of high-fidelity simulations (Appendix F). In addition, MF-(TS)NPE achieved comparable performance with a total computational cost $4.44\pm 0.06$ times lower than standard NPE (Appendix J). Finally, TARP and simulation-based calibration tests suggest that both MF-NPE and NPE estimates were relatively well calibrated (Fig. 3C) (Talts et al., 2020; Lemos et al., 2023).
MF-TSNPE-AF pre-trained on $10^{4}$ low-fidelity samples outperforms MF-NPE trained on $10^{5}$ samples. However, MF-TSNPE-AF performance comes at the cost of training time due to the use of an ensemble of density estimators (Appendix J). This additional training burden is only justified when the simulation cost is substantially higher than the training cost.
Figure 3: (A) Thick-tufted layer 5 pyramidal cell from the neocortex. (B) Performance evaluation with NLTP (same naming convention as in Fig. 2). Amortized methods are averaged over 10 network initializations; non-amortized trained once per 100 observations. Similar results were obtained with NRMSE (Appendix F.1). MF-NPE, and especially its sequential variants, are orders of magnitude more simulation-efficient than NPE. (C) TARP posterior calibration check shows that NPE and MF-NPE trained on $10^{3}$ high-fidelity samples are well-calibrated (Lemos et al., 2023). Simulation-based calibration, posterior samples, and predictives are in Appendix F. (D) Schematic of the low and high-fidelity models of a spiking network. (E) Performance of NPE and MF-NPE evaluated on 10000 true observations with NLTP: averages over 10 network initializations, and 95% confidence intervals. (F) Proportion of posterior samples within the target firing rate bounds. MF-NPE produces a higher fraction of parameter sets within the bounds than NPE.
@@ 4.3 Recurrent spiking network
The high-fidelity simulator consisted of a recurrent network of 4096 excitatory ($E$) and 1024 inhibitory ($I$) leaky integrate-and-fire neurons connected with conductance-based synapses (Fig. 3D). Each synapse type in this network ($E$-to-$E$, $E$-to-II, II-to-$E$, II-to-II) was plastic with an unsupervised local learning rule. For each synapse type, 6 parameters governed how the recent pre- and post-synaptic activity were used to update the synapse, for a total of 24 free parameters across all 4 synapse types (Confavreux et al., 2023). The networks were simulated using Auryn, a C++ simulator (Zenke and Gerstner, 2014) (details in Appendix G).
@@
Summary statistics of the low- and high-fidelity models were the average firing rates of the excitatory and inhibitory neurons at steady state (after 2 minutes of simulation in the high-fidelity model). Plastic networks were considered plausible if the firing rates were between 1 and 50Hz (Dayan and Abbott, 2001; Confavreux et al., 2023).
In this task, the low-fidelity model focuses solely on the $E$-to-$E$ and $E$-to-II rules from the high-fidelity model, thereby having 12 out of the 24 parameters of the high-fidelity model. This setup allows us to demonstrate the performance of MF-NPE on problems with different parameter spaces, highlighting MF-NPE’s flexibility and advantages. We found that MF-NPE has better performance than NPE in terms of NLTP (Fig. 3E), although we observed a diminishing performance gain with increasing discrepancy between the number of parameters of the low- and high-fidelity models (see Appendix G.3). Furthermore, MF-NPE leads to an increase of almost $30\%$ in the proportion of posterior samples within the target firing rate bounds (Fig. 3F), reinforcing that MF-NPE is a practical and effective method for SBI of costly real-world simulators.
Figure 4: (A) Schematic figure representing lower bound on transfer error ($1/\text{MF-NPE{} performance}$) as a function of mutual information between the low- and high-fidelity models, given a fixed simulation budget. (B) Uncertainty coefficient monotonically decreases with noise parameter $\delta$ and is invariant to data inversion. (C) Empirical results with MF-NPE support the hypothesis that transfer performance is dependent on both mutual information and representational coherence. Note that NPE (with the same high-fidelity simulation budget of $10^{2}$) has similar performance as MF-NPE in the case where the low- and high-fidelity models have low mutual information.